Evaluate 2𝑥^3 - 5𝑥^2 + 5𝑥 − 7 at 𝑥 = −3
Answer by ikleyn(52781)

.
Evaluate 2𝑥^3 - 5𝑥^2 + 5𝑥 − 7 at 𝑥 = −3
~~~~~~~~~~~~~~~~

Evaluate, in this case means "substitute x= 3 into the formula and calculate"

    2*(-3)^3 - 5*(-3^2) - 5*(-3) - 7 = 2*(-27) - 5*9 + 15 - 7 = -91.    ANSWER

Solved.

-----------

In this form, evaluating expressions are the easiest problems in Algebra.